### v1

# On the conditioning of histopolationL. Bruni Bruno, S. Serra-Capizzano

Abstract
Histopolation is the approximation procedure that associates a degree d1d-1 polynomial $p_{d-1}\in\mathscr{P}_{d-1}(I)$ with a locally integrable function $f$ imposing that the integral (or, equivalently, the average) of $p$ coincides with that of $f$ on a collection of $d$ distinct segments $s_{i}$. In this work we discuss unisolvence and conditioning of the associated matrices, in an asymptotic linear algebra perspective, i.e., when the matrix-size $d$ tends to infinity. While the unisolvence is a rather sparse topic, the conditioning in the unisolvent setting has a uniform behavior: as for the case of standard Vandermonde matrix-sequences with real nodes, the conditioning is inherently exponential as a function of $d$ when the monomial basis is chosen. In contrast, for an appropriate selection of supports, the Chebyshev basis of second kind has a bounded conditioning. A linear behavior is also observed in the Frobenius norm.
keywords:
Polynomial histopolation; conditioning; Chebyshev polynomials.
{AMS}
15A12, 41A10, 42A15, 65D05
1 Introduction
Let $I\subset\mathbb{R}$ be an interval and let $f:I\to\mathbb{R}$ be a locally integrable function. Histopolation is the approximation procedure that associates with $f$ a degree d1d-1 polynomial $p_{d-1}\in\mathscr{P}_{d-1}(I)$ such that
(1) $\int_{\alpha_{i}}^{\beta_{i}}f(x)\mathrm{d}x=\int_{\alpha_{i}}^{\beta_{i}}p(x)\mathrm{d}x,$
for an appropriate collection of supports $s_{i}:=[\alpha_{i},\beta_{i}]$, $i=1,\ldots,d$. In the literature [5, 21], an equivalent formulation of (1) containing normalization by the measure $|s_{i}|$ of $s_{i}$ is sometimes adopted; in what follows, we stick with the non-normalized case. The main consequence of this choice is that the measures of the supports must be bounded away from zero; that is, we assume that there exists $\varepsilon>0$ independent of $d$ such that $|s_{i}|\geq\varepsilon$ for each $i=1,\ldots,d$. This is a rather natural assumption: even in the normalized case, it avoids the introduction of technical concepts such as regular sets [18]. We will also discuss limiting situations where segments collapse to points.
In contrast with the nodal interpolation procedure, where any collection of $d$ distinct points is unisolvent for polynomial interpolation in $\mathscr{P}_{d-1}(I)$, such a neat recipe does not exist for histopolation, where the identification of specific classes is a hard task. It is indeed very easy to construct examples of $d$ distinct segments that are not unisolvent for histopolation in $\mathscr{P}_{d-1}(I)$.
Example.
Consider the family of segments [i,i][-i,i], for $i=1,\ldots,d$ with $d\geq 1$. Any non-vanishing linear combination of polynomials of odd degree less or equal than d1d-1 satisfies
$\int_{-i}^{i}q_{d-1}(x)\mathrm{d}x=0$
for $i=1,\ldots,d$.
This is only one of the many differences between nodal interpolation and histopolation. Such a dissimilarity is mostly motivated by the fact that any interval $s_{i}$ requires two parameters to be described (e.g., its endpoints $\alpha_{i}$ and $\beta_{i}$, or the midpoint $\tau_{i}$ and the length $\rho$), giving rise to a generally underdetermined problem. As a direct consequence of this, unisolvence results are confined to specific classes of segments. Many of them were identified in [4], and are:
  • (C1)
    chains of intervals. For $i=2,\ldots,d$, one has $\alpha_{i}=\beta_{i-1}$; when $\beta_{i}-\alpha_{i}=|s_{i}|$ is constant for $i=1,\ldots,d$, we call such a collection equidistributed;
[…]
    Figure 1: The segments $s_{1}$, $s_{2}$ and $s_{3}$ form a chain of intervals. They are examples of segments in the class (C1).
  • (C2)
    segments with uniform arc-length. For some $\tau_{i}\in(0,\pi/2)$ and $0<\rho\leq\pi/2$, one has $\alpha_{i}=\cos(\tau_{i}+\rho)$ and $\beta_{i}=\cos(\tau_{i}-\rho)$;
[…]
    Figure 2: The segments $s_{1}$ and $s_{2}$ are the projections of arcs on the semicircle with constant arc-length $2\rho$. They are examples of segments in the class (C2).
  • (C3)
    segments with identical (left or right) endpoint. For $i=1,\ldots,d$, one has $\alpha_{i}=\alpha$, for some $\alpha\in I$, and $\beta_{i}\neq\beta_{j}$ for $i\neq j$. In the right endpoint case, for $i=1,\ldots,d$ one has $\alpha_{i}\neq\alpha_{j}$ for $i\neq j$ and $\beta_{i}=\beta$ for some $\beta\in I$.
[…]
    Figure 3: The segments $s_{1}$, $s_{2}$ and $s_{3}$ share the same left endpoint. They are examples of segments in the class (C3).
Of course, there are families of segments that belong to more than one class. An example of this are the so-called Chebyshev segments, that we will study in depth in Section 3. Indeed, for a particular choice of the arc-length parameter $\rho$, they can be framed both in class (C1) and class (C2). By exploiting the normalized version of (1), segments solving the Fekete problem have subsequently been recognized as an additional scenario [5]. In Section 2.3, we exhibit a new class of segments which is unisolvent for polynomial histopolation. We will denote this class by (C4).
The difficulty in identifying unisolvent classes for polynomial histopolation is reflected in the lack of explicit expressions for the corresponding Lagrange bases, which are known only for the aforementioned classes (C1) [14, 18] and (C3) [4]. The absence of closed formulas for Lagrange polynomials forces to deal with solving a linear system that involves the Vandermonde matrix $V_{d}$, which is hence computed with respect to any other basis $b_{1},\ldots,b_{d}$ of $\mathscr{P}_{d-1}(I)$:
(2) $\left[V_{d}\right]_{i,j}:=\int_{s_{i}}b_{j}(x)\mathrm{d}x.$
As many of our results are asymptotic or directly related with the number of segments, denoted by $d$, we will consistently index the Vandermonde matrix $V_{d}$. This quantity is also the size of the matrix, and is upshifted by $1$ with respect to the relative polynomial degree.
Example.
The non-unisolvence of the set of segments introduced in the above example is also readily seen by looking at the Vandermonde matrix computed with respect the monomial basis $b_{j}(x)=x^{j-1}$. For any even $j$, one has
$\left[V_{d}\right]_{i,j}=\int_{-i}^{i}x^{j-1}\mathrm{d}x=\frac{1}{j}\left[x^{j}\right]_{-i}^{i}=\frac{1}{j}\left(i^{j}-i^{j}\right)=0.$
It follows that any even row of $V_{d}$ is made out of zeros. Since the invertibility of the Vandermonde matrix does not depend on the polynomial basis chosen, one deduces that such a set is not unisolvent for histopolation in $\mathscr{P}_{d-1}(I)$.
The problem of computing the features of a Vandermonde matrix, in particular concerning its conditioning and hence the identification of ill-conditioned bases, is a cornerstone of approximation theory [20]; of course, so is also the opposite problem. In the framework of nodal interpolation, the problem was extensively studied, and several results concerning optimality [12, 13, 16], bad conditioning [2, 6], as well as their applications [17], have been obtained. It is worth pointing out that the majority of these results can be grouped into two clusters: results concerning the monomial basis, which in general offers unstable results, and results concerning the Chebyshev (of first kind) basis, which can be used in combination with Chebyshev nodes to reverse such a situation.
Outline of the paper and main results. In contrast with nodal interpolation, very few is known about the spectral features of the Vandermonde matrix (2) associated with the histopolation problem. The principal aim of this work is to fill this gap.
To do so, we dedicate the next section to ill-conditioning. We recognize that the (Euclidean) conditioning $\kappa_{2}$ of the matrix $V_{d}$ grows exponentially with the variable $d$ when the monomial basis is chosen. This result is presented in Section 2, first for the reference inverval I=[1,1]I=[-1,1] and then extended to any collection of segments by Theorem 2.13. This feature, which recalls the nodal case, is not a consequence of the mean value theorem. To broaden the range of applicability of the result, we also identify a new class of unisolvent segments, and use it to produce numerical evidences.
To restore hope in histopolation, in Section 3 we present results concerning well-conditioning. Indeed, we show that Chebyshev polynomials of the second kind, in combination with segments in the class (C2), give rise to a column-orthogonal matrix. The diagonal terms of this matrix are explicitly determined in Proposition 3.1. As a consequence, in Theorem 3.8 we prove that the (Euclidean) conditioning $\kappa_{2}(V_{d})$ is bounded, and in Theorem 3.14 that the (Frobenius) conditioning $\kappa_{F}(V_{d})$ is linear. Both these results are provided with simple formulas that depend on $\rho$.
2 Ill-conditioning: the monomial basis
The principal aim of this section is to prove that the Euclidean conditioning
$\kappa_{2}\left(V_{d}\right)=\frac{\sigma_{1}\left(V_{d}\right)}{\sigma_{d}\left(V_{d}\right)}=\sqrt{\frac{\lambda_{1}\left(V_{d}^{\top}V_{d}\right)}{\lambda_{d}\left(V_{d}^{\top}V_{d}\right)}}$
of the Vandermonde matrix $V_{d}$ defined in (2) is exponential when the monomial basis $b_{j}(x):=x^{j-1}$ is chosen. As we shall see, this holds independently of the underlying collection of unisolvent segments. We build such a general result by first consdering the peculiar case of the class (C3), which yields very neat computations, and then show how our proof extends to other scenarios. To fix the ideas we start with a motivating example, whose features turn out to be quite general.
2.1 A motivating example
Consider $s_{i}=[0,i]$, for $i=1,\ldots,d$ and the monomial basis. The corresponding Vandermonde matrix is
$\left[V_{d}\right]_{i,j}=\int_{0}^{i}x^{j-1}\mathrm{d}x=\frac{i^{j}}{j},$
for which one readily computes that
$\left[V_{d}^{\top}V_{d}\right]_{i,j}=\frac{1}{i\cdot j}\sum_{k=1}^{d}k^{i+j}=:\left[W_{d}\right]_{i,j}.$
By the well-known properties of the Raylegh quotient, for any $x\in\mathbb{R}^{d}$ one has
$\lambda_{d}\left(W_{d}\right)\leq\frac{x^{\top}W_{d}\,x}{x^{\top}x}\leq\lambda_{1}\left(W_{d}\right).$
By considering $x=e_{1}$, the first element of the canonical basis, one immediately gets that
$\lambda_{d}\left(W_{d}\right)\leq\left[W_{d}\right]_{1,1}=\sum_{k=1}^{d}k^{2}=\frac{d(d+1)(2d+1)}{6}$
and likewise, taking $x=e_{d}$, one also has that
$\lambda_{1}\left(W_{d}\right)\geq\left[W_{d}\right]_{d,d}=\frac{1}{d^{2}}\sum_{k=1}^{d}k^{2d}>d^{2d-2}.$
It follows that, for this specific case, the conditioning of $V_{d}$ is exponential with respect to the polynomial degree $d$. By the linearity of the integral, a very similar result can be obtained for the case (C1) when segments $s_{i}=[i-1,i]$ are considered. These facts are pictured in Figure 4. In what follows, we will sharpen these estimates, and remove the specific assumption on the supports.
Figure 4: Exponential conditioning of the Vandermonde matrix in the monomial basis for classes (C3) and (C1), left and right hand panel respectively.
[…]
The motivating example is important for the following reason. Suppose that there exists $c>0$ such that $\sup_{i}\left(\beta_{i}-\alpha_{i}\right)\geq c$. When there exists $v>1$ independent of $d$ such that $\sup_{i,j}\max\{|\beta_{i}|,|\alpha_{j}|\}\geq v$, by direct inspection as above we have necessarily exponential growth for the maximal eigenvalue of $W_{d}=V_{d}^{\top}V_{d}$, with its minimal eigenvalue that cannot grow exponentially with respect to $d$. Hence, the latter gives a clear evidence of an exponential conditioning of $W_{d}$ and a fortiori of $V_{d}$ as a function of the size $d$, whenever $c>0$, $v>1$ are independent of the polynomial degree $d$. In what follows, we shall thus make a similar regularity assumption on $c$.
Now, observe that dividing by the size $d$ of the longest support $s_{d}$ one translates the problem in a fixed reference interval. By applying the change of variable $t=x/d$, one immediately sees that the $j$-th column of $V_{d}$ scales by the factor $1/d^{j}$ and, as a consequence, the $(i,j)$-th term of $W_{d}$ scales by $1/d^{i+j}$. This shows that the conditioning of $V_{d}$ depends on the length of the segments. In particular, even if it is easy to numerically check that the conditioning in this situation is still exponential, this example complicates the estimate on $\lambda_{1}$ given above and suggest that the reference interval I=[1,1]I=[-1,1] is somewhat a limiting case of the problem. This will guide our proof strategy.
2.1.1 A formula for the determinant for class (C3)
Before moving further, we extend [5, Eq. (24)], and obtain an easy formula for the determinant of $V_{d}$ for the class (C3) when the left endpoint is $\alpha_{i}=\alpha$. The special case $\alpha=0$ clearly applies to the motivating example.
Lemma 2.1.
Let $s_{i}=[\alpha,\beta_{i}]$, with $\alpha<\beta_{i}<\beta_{j}$ for $1\leq i<j\leq d$. The determinant of $V_{d}$ is positive, and
(3) $\det V_{d}=\frac{1}{d!}\left(\prod_{i=1}^{d}\left(\beta_{i}-\alpha\right)\right)\left(\prod_{1\leq j<k\leq d}\left(\beta_{k}-\beta_{j}\right)\right).$
Proof 2.2.
We establish formula (3), the claim on the positivity will be a trivial consequence of the ordering $\alpha<\beta_{i}<\beta_{j}$ for $1\leq i<j\leq d$. We compute
$V_{d}]_{i,j}=\int_{\alpha}^{\beta_{i}}x^{j-1}\mathrm{d}x=\frac{1}{j}\left(\beta_{i}^{j}-\alpha^{j}\right),$
with $j=1,\ldots,d$. Then $V_{d}$ factorizes as $U_{d}D_{d}$, where $D_{d}$ is the diagonal matrix $D_{d}]_{j,j}=j^{-1}$ and $U_{d}$ is the shifted Vandermonde matrix whose $(i,j)$-th element is
$U_{d}]_{i,j}=(\beta_{i}^{j}-\alpha^{j})=(\beta_{i}-\alpha)\sum_{k=0}^{j-1}\alpha^{k}\beta_{i}^{j-1-k}.$
Gathering the term $(\beta_{i}-\alpha)$ row-wise in $U_{d}$, we may write $U_{d}=B_{d}M_{d}$, where $B_{d}$ is the diagonal matrix with entries $B_{d}]_{i,i}=(\beta_{i}-\alpha)$ and
$\left[M_{d}\right]_{i,j}:=\sum_{k=0}^{j-1}\alpha^{k}\beta_{i}^{j-1-k}.$
By subtracting $\alpha$ times the $j$-th column from the $(j+1)$-th in $M_{d}$, for $j=d-1,\ldots,1$, one finds that
$\det M_{d}=\det\left(\left[\beta_{i}^{j-1}\right]_{i,j=1}^{d}\right)$
The matrix appearing at the right hand side of the above equation is the Vandermonde matrix associated with the nodes $\{\beta_{1},\ldots,\beta_{d}\}$, whose determinant is well-known to be $\prod_{1\leq j<k\leq d}\left(\beta_{k}-\beta_{j}\right)$. The claim immediately now follows from Binet’s Theorem.
The formula (3) assumes the very neat form $\det(V_{d})=\prod_{i=1}^{d-1}i!$ under the assumptions of Section 2.1. This is readily seen plugging $\alpha=0$ and $\beta_{i}=i$ in (3). Lemma 2.1 also gives an alternative, more direct proof of the unisolvence of the class (C3) proposed in [4].
2.1.2 An application to the Fekete problem
The Fekete problem has been celebrated in interpolation theory since its definition [11]. Solutions to nodal Fekete problems emerge in uniform approximation, as the corresponding Lagrange bases have small norm. As a consequence, the relative Lebesgue constant is small, offering good numerical stability. This holds for the segmental Fekete problem as well [5], provided that one considers the normalized Fekete problem: find (possibly degenerate) segments $\{s_{1},\ldots,s_{d}\}$ such that
(4) $\left|\det\left(\left[\frac{1}{|s_{i}|}\int_{s_{i}}b_{j}(x)\mathrm{d}x\right]_{i,j}\right)\right|$
is maximized. Clearly, such a procedure does not depend on the polynomial basis chosen [3], and this fact is extensively used in [5] to obtain explicit solutions for the classes (C1) and (C2). Owing to this and (3), we are in a position to provide a solution for the class (C3).
Proposition 2.3.
Fix $\alpha\in I$, and let $s_{i}:=[\alpha,\beta_{i}]$. Segments $\{s_{1},\ldots,s_{d}\}$ that maximize (4) are such that $\{\beta_{1},\ldots,\beta_{d}\}$ are the Fekete nodes of degree $d$.
Proof 2.4.
For notational simplicity, let us assume that $\alpha<\beta_{1}<\ldots<\beta_{d}$. Call $\widetilde{V}_{d}$ the normalized Vandermonde matrix of (4), and observe that
$\det\widetilde{V}_{d}=\prod_{i=1}^{d}\frac{1}{|s_{i}|}\cdot\det V_{d}=\frac{1}{\prod_{i=1}^{d}\left(\beta_{i}-\alpha\right)}\det V_{d}.$
Consider the monomial basis $b_{j}(x)=x^{j-1}$. By (3), one has
$\det\widetilde{V}_{d}$ $=\frac{1}{\prod_{i=1}^{d}\left(\beta_{i}-\alpha\right)}\det V_{d}=\frac{1}{\prod_{i=1}^{d}\left(\beta_{i}-\alpha\right)}\frac{1}{d!}\left(\prod_{i=1}^{d}\left(\beta_{i}-\alpha\right)\right)\left(\prod_{1\leq j<k\leq d}\left(\beta_{k}-\beta_{j}\right)\right)$
$=\frac{1}{d!}\prod_{1\leq j<k\leq d}\left(\beta_{k}-\beta_{j}\right),$
where we also removed the absolute value of (4) due to the positivity result given in Lemma 2.1. Since $\frac{1}{d!}$ is a constant term, $\det\widetilde{V}_{d}$ is maximixed by nodes $\{\beta_{1},\ldots,\beta_{d}\}$ that maximize the rightmost term of the above equality. Such nodes are well-known to be the Fekete points, which is the claim. Notice that the case $\alpha=\beta_{1}$ also follows by an immediate continuity argument.
2.2 A conditioning result for the reference interval
To extend the computations of Section 2.1, we start from the situation where the reference interval is I=[1,1]I=[-1,1]. This framework offers a neat handling due to the well-known features of Chebyshev polynomials.
Theorem 2.5.
Let $1\leq\alpha_{i}<\beta_{i}\leq 1$ for $i=1,\ldots,d$, and assume that the corresponding collection of segments is unisolvent for histopolation in $\mathscr{P}_{d-1}(I)$. Suppose that there exists $c$ independent of $d$ such that $\max_{i}\left(\beta_{i}-\alpha_{i}\right)\geq c>0$. Then
$\kappa_{2}\left(V_{d}\right)\geq\frac{c}{\sqrt{d}}2^{d-2}.$
Proof 2.6.
By (2), we immediately get that
$\left[V_{d}\right]_{i,j}=\frac{\beta_{i}^{j}-\alpha_{i}^{j}}{j}.$
Let $e_{i}$ be the $i$-the element of the canonical basis of $\mathbb{R}^{d}$. Then, lower bounding the $2$-norm of the vector $V_{d}e_{i}$ by its first component, one has
$\left\|V_{d}e_{i}\right\|_{2}=\left\|\left[\frac{\beta_{i}^{j}-\alpha_{i}^{j}}{j}\right]_{j=1}^{d}\right\|_{2}\geq\left|\beta_{i}-\alpha_{i}\right|=\beta_{i}-\alpha_{i},$
whence
$\left\|V_{d}\right\|_{2}=\max_{\|u\|_{2}=1}\left\|V_{d}u\right\|_{2}\geq\max_{1\leq i\leq d}\left\|V_{d}e_{i}\right\|_{2}\geq\max_{1\leq i\leq d}\beta_{i}-\alpha_{i}\geq c,$
for a $c$ independent of the polynomial degree (hence of the matrix size) $d$. Now, in order to estimate from below the conditioning of $V_{d}$ in the Euclidean norm, we look for a proper estimation from below of $\left\|V_{d}^{-1}\right\|_{2}$. In particular, observe that for any $y\in\mathbb{R}^{d}$ one has
(5) $\left\|V_{d}^{-1}\right\|_{2}=\max_{\begin{subarray}{c}z\in\mathbb{R}^{d}\\ z\neq 0\end{subarray}}\frac{\|V_{d}^{-1}z\|_{2}}{\|z\|_{2}}\geq\frac{\|y\|_{2}}{\left\|V_{d}y\right\|_{2}}.$
It follows that, if we are able to select an appropriate $y\in\mathbb{R}^{d}$, we obtain a satisfactory estimate for $\left\|V_{d}^{-1}\right\|_{2}$. To this end, consider the Chebyshev polynomial $p_{d-1}^{*}(x):=\sum_{j=1}^{d}a_{j}x^{j-1}$ of degree d1d-1. This is a monic polynomial, so $a_{d}=1$, and it is well-known that its norm on [1,1][-1,1] is $\|p_{d-1}^{*}\|_{\infty}=2^{1-d}$. We evaluate the inverse of $V_{d}$ on the transformed vector $b=V_{d}a$. By construction, we have
$b_{i}=\int_{\alpha_{i}}^{\beta_{i}}p_{d-1}^{*}(z)\mathrm{d}z,$
whence, by the Hölder inequality,
$|b_{i}|=\left|\int_{\alpha_{i}}^{\beta_{i}}p_{d-1}^{*}(z)\mathrm{d}z\right|\leq(\beta_{i}-\alpha_{i})\|p_{d-1}^{*}\|_{\infty}\leq(\beta_{i}-\alpha_{i})2^{1-d}.$
Therefore
$\|b\|_{2}\leq\max_{1\leq i\leq d}(\beta_{i}-\alpha_{i})\frac{\sqrt{d}}{2^{d-1}}\leq 2\frac{\sqrt{d}}{2^{d-1}}=\frac{\sqrt{d}}{2^{d-2}}.$
By (5) we immediately deduce that $\left\|V_{d}^{-1}\right\|_{2}\geq 1/\|b\|_{2}\geq 2^{d-2}/\sqrt{d}$. Gathering all the results, we have
$\kappa_{2}\left(V_{d}\right)=\|V_{d}\|_{2}\|V_{d}^{-1}\|_{2}\geq\frac{c}{\sqrt{d}}2^{d-2}.$
Since $c$ does not depend on the degree $d$, the claim is proved.
Theorem 2.5 applies to the class of segments (C1) where $1<\beta_{i}=\alpha_{i+1}<1$ for $i=1,\ldots,d-1$. It contains also the class of segments (C3) for $1\leq\alpha=\alpha_{i}$ for $i=1,\ldots,d$ and $\beta_{i}\neq\beta_{j}$ for $i\neq j$. Even if confined to the reference element I=[1,1]I=[-1,1], Theorem 2.5 proves that the trend depicted in Figure 4 is not peculiar to the selection of segments performed in the motivating Section 2.1. The subsequent Theorem 2.13 will in fact demonstrate that this is intrinsic to the monomial basis.
2.3 Segments with uniform length: the class (C4)
The technical assumption $\max_{i}(\beta_{i}-\alpha_{i})\geq c>0$ makes the result of Theorem 2.5 tailored for a new class of segments, in which the constant $c$ is the measure of a given reference element.
Definition 2.7.
Let $s=[\alpha,\beta]$ be a reference segment and let $\{\xi_{1},\ldots,\xi_{d}\}$ be real nodes. We define the segments in the class (C4) as the collection $\{s_{1},\ldots,s_{d}\}$, where
$s_{i}:=\{x+\xi_{i}:\ x\in s\}.$
Clearly, one has $s_{i}=[\alpha+\xi_{i},\beta+\xi_{i}]$ and $|s_{i}|=|s|$ for each $i=1,\ldots,d$.
[…]
Figure 5: The segments $s_{1}$, $s_{2}$ and $s_{3}$ are translated of the same segment. They are examples of segments in the class (C4).
The class (C4) shows that the way segments overlap is a rather weak information for the purpose of unisolvence. In fact, we are going to show that the class (C4) is unisolvent for histopolation in $\mathscr{P}_{d-1}(I)$ is unisolvent whenever the nodes $\xi_{i}$, $i=1,\ldots,d$, are pairwise distinct, with no restriction on the measure of $s$. This can be seen as a consequence of the following general fact.
Lemma 2.8.
Let $f:\mathbb{R}\to\mathbb{R}$ be a continuous function, and let $s$ be an interval. Define $s+\xi:=\{x+\xi:\ x\in s\}$. If
$\int_{s+\xi}f(x)\mathrm{d}x=0$
for each $x\in\mathbb{R}$, then $f(x)=0$ for each $x\in\mathbb{R}$.
Proof 2.9.
For ease, observe that the result can be restated in terms of translations by $\xi$ in place of $\xi$ and consider the characteristic function $\chi_{s-\xi}$ for the interval $s-\xi$. Thus, for each $\xi\in\mathbb{R}$, one has
$0=\int_{s-\xi}f(x)\mathrm{d}x=\int_{\mathbb{R}}\left(f\cdot\chi_{s-\xi}\right)(x)\mathrm{d}x=\left(f\ast\chi_{s-\xi}\right)(x),$
begin $\ast$ the usual convolution operator. By construction, one has that $f\cdot\chi_{s-\xi}(x)\in L^{1}(\mathbb{R})$, and we may consider the action of the Fourier transform $\mathcal{F}$ between integrable functions in $L^{1}$ and bounded functions in $L^{\infty}$, which gives
(6) $0=\mathcal{F}(0)=\mathcal{F}(f\ast\chi_{s-\xi})=\mathcal{F}(f)\cdot\mathcal{F}(\chi_{s-\xi}).$
By direct computation, setting $s=[\alpha,\beta]$, one obtains
$\left(\mathcal{F}(\chi_{s-\xi})\right)(t)=e^{i\pi t\left(2\xi-(\alpha+\beta)\right)}\frac{\sin\left(|s|t\right)}{t},$
continuously extended at $t=0$. The norm of this function is bounded by the measure of $s$, since $|\mathcal{F}(\chi_{s-\xi})|\leq|s-\xi|=|s|<+\infty$. Further, for each $\xi$ the function $\mathcal{F}(\chi_{s-\xi})$ vanishes on a discrete set. By (6), it must be $\mathcal{F}(f)=0$ almost everywhere, and the continuity of $f$ implies that $f=0$.
We are in a position to prove the following unisolvence result.
Proposition 2.10.
Let $s$ be a segment, and define $s+\xi:=\{x+\xi:\ x\in s\}$. Consider $\left\{\xi_{1},\ldots,\xi_{d}\right\}$ pairwise distinct nodes and let $p_{d-1}\in\mathscr{P}_{d-1}(I)$. Then, if
$\int_{s+\xi_{i}}p_{d-1}(x)\mathrm{d}x=0$
for each $i=1,\ldots,d$, then $p_{d-1}(x)=0$ for each $x\in\mathbb{R}$.
Proof 2.11.
Consider the map
$\xi\mapsto\int_{s+\xi}p_{d-1}(x)\mathrm{d}x.$
By the change of variable, this is a polynomial map (in the unknown $\xi$) of degree d1d-1. By hypothesis, this polynomial vanishes on the grid $\left\{\xi_{1},\ldots,\xi_{d}\right\}$, which consists of distinct nodes. Hence this polynomial vanishes on the whole real line $\mathbb{R}$. By Lemma 2.8, one gets $p_{d-1}(x)=0$ for each $x\in\mathbb{R}$.
Remark 2.12.
Proposition 2.10 also has a different proof. To this end, observe that the set of zeros of $p_{d-1}(x)$ is a discrete set, containing at most d1d-1 nodes. If it had more, then $p_{d-1}$ would be the zero polynomial. Call $\xi_{\max}$ the largest one, i.e. the node $\xi_{i}$ such that $\xi_{i}\geq\xi_{j}$ for any $i\neq j$. Then, the result of Lemma 2.8 can be deduced in the specific case of polynomials as there exists $\xi\in\mathbb{R}$ such that $\xi-|s|\geq\xi_{\max}$. Hence, the sign of $p_{d-1}$ is constant in the interior of $s+\xi$, and since $\int_{s+\xi}p_{d-1}(x)\mathrm{d}x=0$, one has $p_{d-1}(x)=0$ for each $x$ in the (non empty) interior of $s+\xi$. This implies that $p_{d-1}(x)=0$ for each $x\in\mathbb{R}$.
The class (C4) has intersection with the class (C1), which is given by what we called, in the introduction, equidistributed segments. Due to Theorem 2.5, The Vandermonde matrix associated with such segments and the monomial basis presents asymptotically an exponential ill-conditioning as a function of $d$. This is numerically verified and shown in Figure 6.
Figure 6: Exponential asymptotic ill-conditioning for segments in the class (C4).
2.4 Extension to generic intervals
The extension of Theorem 2.5 to the generic case is based on the following result.
Theorem 2.13.
Let $\alpha_{i}<\beta_{i}$ for $i=1,\ldots,d$ and assume $\liminf_{d\rightarrow\infty}\min_{i}\alpha_{i}=-1$ or, equivalently, $\limsup_{d\rightarrow\infty}\max_{i}\beta_{i}=1$. Suppose that there exists $c$ independent of $d$ such that $\max_{i}(\beta_{i}-\alpha_{i})\geq c>0$. Then
$\kappa_{2}\left(V_{d}\right)\geq\frac{c}{\sqrt{d}}2^{d-2}.$
Proof 2.14.
We assume $\limsup_{d\rightarrow\infty}\max_{i}\beta_{i}=1$ and let $i^{*}_{d}$ such that $\beta_{i^{*}_{d}}=\max_{1\leq i\leq d}\beta_{i}$. With no loss of generality we can assume $\alpha_{i}\geq-1$. Hence
$\frac{\beta_{i}^{j}-\alpha_{i}^{j}}{j}=\beta_{i^{*}_{d}}^{j}\frac{\gamma_{i}^{j}-\delta_{i}^{j}}{j},$
[…]
$\kappa_{2}(\widehat{V}_{d})\leq\kappa_{2}(V_{d})\kappa_{2}(D_{d})$
i.e.
$\kappa_{2}(V_{d})\geq\frac{\kappa_{2}(\widehat{V}_{d})}{\kappa_{2}(D_{d})}.$
Now, taking into account the hypothesis $\limsup_{d\rightarrow\infty}\max_{i}\beta_{i}=1$, we infer that for each $\varepsilon>0$, there exists $d_{\varepsilon}$ such $\kappa_{2}(D_{d})\leq(1+\varepsilon)^{d}$, for any $d\geq d_{\varepsilon}$ and finally by Theorem 2.5 we conclude the proof since
$\kappa_{2}(V_{d})\geq\frac{\kappa_{2}(\widehat{V}_{d})}{\kappa_{2}(D_{d})}\geq\frac{c}{\sqrt{d}}\frac{2^{d-2}}{(1+\varepsilon)^{d}}$
and $\varepsilon$ is chosen so that $i+\varepsilon<2$.
To better frame Theorem 2.13, we conclude the section gathering some considerations and consequences.
Remark 2.15.
In connection with the proof of Theorem 2.13, we note that the latter is the missing piece of the picture. Indeed, we are finally in the position of completing the reasoning with the following observation: if $\liminf_{d\rightarrow\infty}\lim_{i}\alpha_{i}<-1$ or $\limsup_{d\rightarrow\infty}\max_{i}\beta_{i}>1$, then passing to proper subsequences, we go back to the same setting as in Remark 2.1. As a conclusion the exponential conditioning is a general phenomenon. This is particularly interesting given the link of the considered structures with the Vandermonde matrices on with nodes on the real line; see [2, 19].
Remark 2.16.
If we consider standard induced norms $\|\cdot\|_{l^{k}\rightarrow l^{p}}$, $k,p\in[1,\infty]$ (the case $k=p=2$ being the standard Euclidean case, also known as spectral norm), then we obtain again exponential conditioning, since the equivalence constants among all these induced norms have a rational expression with respect to the variable $d$.
Remark 2.17.
The picture given in Theorems 2.5 and 2.13 and in Remarks 2.1 and 2.15 is even more general. In fact, if we consider sets $\{i\cdot\alpha_{i}^{(d)}=1,\ldots,d\}$, $\{i\cdot\beta_{i}^{(d)}=1,\ldots,d\}$, of natural numbers, in which for $d_{1}\neq d_{2}$ we have in general $\alpha_{i}^{(d_{1})}\neq\alpha_{i}^{(d_{2})}$ and $\beta_{i}^{(d_{1})}\neq\beta_{i}^{(d_{2})}$, $i\leq\min\{d_{1},d_{2}\}$, all the proofs work unchanged. This represents a further degree of generality in the analysis.
2.5 Some remarks on the independence of $c$ on $d$
We have so far assumed the existence of a “separation distance” $c>0$ such that $\max_{i}\left(\beta_{i}-\alpha_{i}\right)\geq c$. We asked such a constant to be independent of $d$. This is clearly motivated by the fact that, if $c$ depended on $d$, it could decay to zero fast enough to simplify the exponential term $2^{d-2}$. It is plain to observe that, if $c$ depends on $d$ and does not decay exponentially, the exponential trend of the conditioning of the monomial basis remains unchanged. An example is provided when considering equidistributed segments in I=[1,1]I=[-1,1].
Suppose, in complete contrast, that such a constant $c=\max_{i}(\beta_{i}-\alpha_{i})$ is equal to zero; this means that every segment $s_{i}$ collapses to a point $\xi_{i}$. Even if these nodes are distinct, the problem (1) becomes ill-posed, as the matrix (2) becomes the zero matrix. This situation can be resolved by placing the normalization factor $|s_{i}|$ in the histopolation conditions, as discuessed in [5]. An immediate continuity argument, together with the mean value theorem, shows that the Vandermonde matrix (2) converges (in an entry-wise sense) to the nodal Vandermonde matrix based upon the nodes $\{\xi_{1},\ldots,\xi_{d}\}$. The well-established nodal theory [2, 19] hence ensures that Theorem 2.13 holds even in this limit.
3 Well-conditioning: Chebyshev basis and segments
In stark contrast with Section 2, a very different situation appears when the Chebyshev polynomials of the second kind
$U_{j}(x)=\frac{\sin(j\arccos x)}{\sin(\arccos x)}$
are considered. This family is not the most diffused in nodal interpolation, where the first one
$T_{j}(x)=\cos(j\arccos x)$
is usually adopted. The two families are related via differentiation, as $\mathrm{d}/\mathrm{d}x\left(T_{j}(x)\right)=jU_{j-1}(x)$. The Vandermonde matrix (2) assumes a very peculiar form with respect to the basis $U_{j}(x)$:
(7) $\left[V_{d}\right]_{i,j}$ $=\int_{\alpha_{i}}^{\beta_{i}}U_{j}(x)\mathrm{d}x=\frac{1}{j}\int_{\alpha_{i}}^{\beta_{i}}T_{j+1}^{\prime}(x)\mathrm{d}x$
$=\frac{2}{j}\sin\left(j\frac{\arccos\alpha_{i}+\arccos\beta_{i}}{2}\right)\sin\left(j\frac{\arccos\alpha_{i}-\arccos\beta_{i}}{2}\right).$
We show that this basis is particular well conditioned when considered in combination with segments in the class (C2). Recall that these segments are defined as
(8) $s_{i}=[\alpha_{i},\beta_{i}]=[\cos(\tau_{i}+\rho),\cos(\tau_{i}-\rho)],$
being $\tau_{i}$ the (pre-image of the) $i$-th Chebyshev node $\tau_{i}=\frac{(i+1/2)\pi}{d}$ for $i=0,\ldots,d-1$ and $\rho>0$ an arc length veryfing the existence conditions established in [4, Section 3.2]. Plugging the definition of the polynomial $U_{j}(x)$ and class (C2) segments into (7), one immediately obtains that the respective Vandermonde matrix is
(9) $\left[V_{d}\right]_{i,j}=\frac{2}{j}\sin(j\tau_{i})\sin(j\rho).$
This matrix has orthogonal columns.
Proposition 3.1.
The matrix $V_{d}^{\top}V_{d}$ is diagonal. The $i$-th diagonal term is
(10) $\left[V_{d}^{\top}V_{d}\right]_{i,i}=\frac{2d}{i^{2}}\sin^{2}(i\rho).$
Proof 3.2.
By direct computation, one obtains
$\left[V_{d}^{\top}V_{d}\right]_{i,j}$ $=\sum_{k=1}^{d}\left[V_{d}^{\top}\right]_{i,k}\left[V_{d}\right]_{k,j}=\sum_{k=1}^{d}\left[V_{d}\right]_{k,i}\left[V_{d}\right]_{k,j}==\sum_{k=1}^{d}\frac{2}{i}\sin(i\tau_{k})\sin(i\rho)\frac{2}{j}\sin(j\tau_{k})\sin(j\rho)$
$=\frac{4\sin(i\rho)\sin(j\rho)}{ij}\sum_{k=1}^{d}\sin(i\tau_{k})\sin(j\tau_{k}).$
By Werner’s formula, we rewrite any term in the sum as
$\sin(i\tau_{k})\sin(j\tau_{k})=\frac{1}{2}\left[\cos((i-j)\tau_{k})-\cos((i+j)\tau_{k})\right].$
Consider first the case $i=j$, i.e. the diagonal terms of $V_{d}^{\top}V_{d}$. One has
$\sum_{k=1}^{d}\sin(i\tau_{k})\sin(j\tau_{k})$ $=\frac{1}{2}\sum_{k=1}^{d}\left[\cos((i-i)\tau_{k})-\cos((i+i)\tau_{k})\right]=\frac{1}{2}\sum_{k=1}^{d}\left[\cos(0)-\cos(2i\tau_{k})\right]$
$=\frac{1}{2}\sum_{k=1}^{d}1-\frac{1}{2}\sum_{k=1}^{d}\cos(2i\tau_{k})=\frac{d}{2}-0=\frac{d}{2},$
where the equality $\sum_{k=1}^{d}\cos(2i\tau_{k})=0$ is an easy (and well-known) consequence of the symmetry of the Chebyshev nodes. The symmetry of the Chebyshev nodes further implies that, for $i\neq j$,
$\sum_{k=1}^{d}\cos\left((i-j)\tau_{k}\right)=\sum_{k=1}^{d}\cos\left((i+j)\tau_{k}\right)=0.$
Hence one has
$\left[V_{d}^{\top}V_{d}\right]_{i,j}=\left[\frac{4}{ij}\sin(i\rho)\sin(j\rho)\right]\frac{d}{2}\delta_{i,j}=\left[\frac{2d}{ij}\sin(i\rho)\sin(j\rho)\right]\delta_{i,j},$
which shows the claim.
Proposition 3.1 is the key ingredient to unveil some features of $V_{d}$ and the corresponding matrix sequence. The first easy consequence is an exact formula for the determinant of $V_{d}$. This formula assumes a peculiar form for $\rho=\pi/(2d)$. The relevance of this case is described in [4], and will give further simplifications also for the next sections.
Corollary 3.3.
One has
(11) $\det V_{d}=\sqrt{\frac{(2d)^{d}}{(d!)^{2}}\prod_{i=1}^{d}\sin^{2}\left(i\rho\right)}.$
In particular, if $\rho=\pi/(2d)$, one has
$\det V_{d}=\sqrt{\frac{d(2d)^{d}}{(d!2^{d-1})^{2}}}.$
Proof 3.4.
It is clear that $\det(V_{d}^{\top}V_{d})=\left(\det V_{d}\right)^{2}$, whence
$\det V_{d}=\sqrt{\det(V_{d}^{\top}V_{d})}=\sqrt{\prod_{i=1}^{d}[V_{d}^{\top}V_{d}]_{i,i}}=\sqrt{\prod_{i=1}^{d}\frac{2d}{i^{2}}\sin^{2}(i\rho)}.$
Let us examine the argument of the last square root. One has
(12) $\prod_{i=1}^{d}\frac{2d}{i^{2}}\sin^{2}(i\rho)=\frac{\prod_{i=1}^{d}2d}{\prod_{i=1}^{d}i^{2}}\prod_{i=1}^{d}\sin^{2}(i\rho)=\frac{(2d)^{d}}{(d!)^{2}}\prod_{i=1}^{d}\sin^{2}(i\rho),$
which gives the first claim. For what concerns the specialization to $\rho=\pi/(2d)$, we shall invoke the trigonometric equality
(13) $\prod_{i=1}^{2d-1}\sin\left(\frac{i\pi}{2d}\right)=\frac{2d}{2^{2d-1}}.$
Notice, in particular, that we can write
$\prod_{i=1}^{2d-1}\sin\left(\frac{i\pi}{2d}\right)=\prod_{i=1}^{d-1}\sin\left(\frac{i\pi}{2d}\right)\cdot\sin{\left(\frac{d\pi}{2d}\right)}\cdot\prod_{i=d+1}^{2d-1}\sin\left(\frac{i\pi}{2d}\right)=\prod_{i=1}^{d-1}\sin\left(\frac{i\pi}{2d}\right)\cdot\prod_{i=d+1}^{2d-1}\sin\left(\frac{i\pi}{2d}\right).$
By the symmetry of the sine function $\sin(\pi-t)=\sin(t)$, by the change of indices j=2dij=2d-i, one immediately retrieves that
$\prod_{i=d+1}^{2d-1}\sin\left(\frac{i\pi}{2d}\right)=\prod_{i=1}^{d-1}\sin\left(\frac{i\pi}{2d}\right)=\prod_{i=1}^{d}\sin\left(\frac{i\pi}{2d}\right),$
where the last equality follows as the $d$-th term of the product is $\sin(\pi/2)=1$. Plugging this into (13), one obtains
$\prod_{i=1}^{d-1}\sin^{2}\left(\frac{i\pi}{2d}\right)=\frac{2d}{2^{2d-1}}=\frac{d}{2^{2(d-1)}}.$
Plugging this into (12), one finds
$\prod_{i=1}^{d}\frac{2d}{i^{2}}\sin^{2}(i\rho)=\frac{(2d)^{d}}{(d!)^{2}}\frac{d}{2^{2(d-1)}}=\frac{d(2d)^{d}}{(d!2^{d-1})^{2}},$
which immediately yields the claim by taking the square root.
We close this section with an interesting fact. The identity (13) is known, but the authors struggled to find a reference. In fact, it appears in Wolfram MathWorld as a “personal communication”. For completeness, we report a possible proof.
Lemma 3.5.
The trigonometric equality (13) holds; that is, for any $n\in\mathbb{N}$, one has
$\prod_{i=1}^{n-1}\sin\left(\frac{i\pi}{n}\right)=\frac{n}{2^{n-1}}.$
Proof 3.6.
By multiplying both sides for $2^{n-1}$ and using Euler’s formula, we may rewrite left hand side of the identity as
$\prod_{j=1}^{n-1}2\sin\left(\frac{j\pi}{n}\right)$ $=\prod_{j=1}^{n-1}2\sin\left(\frac{j\pi}{n}\right)=\prod_{j=1}^{n-1}\frac{e^{i\frac{j\pi}{n}}-e^{i\frac{j\pi}{n}}}{i}=(-i)^{n-1}\prod_{j=1}^{n-1}e^{-i\frac{j\pi}{n}}\left(e^{2i\frac{j\pi}{n}}-1\right)$
$=(-i)^{n-1}\prod_{j=1}^{n-1}e^{-i\frac{j\pi}{n}}\prod_{j=1}^{n-1}\left(e^{2i\frac{j\pi}{n}}-1\right)=(-i)^{n-1}e^{-i\sum_{j=1}^{n-1}\frac{j\pi}{n}}\prod_{j=1}^{n-1}\left(e^{2i\frac{j\pi}{n}}-1\right)$
$=(-i)^{n-1}e^{-i(n-1)\pi}\prod_{j=1}^{n-1}\left(e^{2i\frac{j\pi}{n}}-1\right)=\prod_{j=1}^{n-1}\left(1-e^{2i\frac{j\pi}{n}}\right).$
We shall prove that this quantity now equals to $n$. The last term can be treated analyzing the roots of the unity, as
$z^{n}-1=\prod_{j=0}^{n-1}\left(z-e^{2i\frac{j\pi}{n}}\right)=(z-1)\prod_{j=1}^{n-1}\left(z-e^{2i\frac{j\pi}{n}}\right),$
whence, by dividing both sides by z1z-1 and computing the quotient, it follows that
$\prod_{j=1}^{n-1}\left(z-e^{2i\frac{j\pi}{n}}\right)=\sum_{j=0}^{n-1}z^{j}.$
Evaluating the expression at $z=1$ we get the claim.
3.1 An inversion formula for the Chebyshev segments
The orthogonality of the Vandermonde matrix can be used to determine its inverse, and in particular the Lagrange basis associated with segments in the class (C2).
Remark 3.7.
The columns of $V_{d}^{-1}$ are the coefficients that turn a given basis into the Lagrange one, i.e. the only (up to permutations) basis $\{\ell_{1},\ldots,\ell_{d}\}$ for $\mathscr{P}_{d-1}(I)$ such that $\int_{s_{i}}\ell_{j}(x)\mathrm{d}x=\delta_{i,j}$, see [6, Proposition 1]. An explicit formula for the inverse of $V_{d}$ is crucial, as such bases in histopolation are known only for segments in the classes (C1) and (C3).
For ease of the next computations, let us set once more $W_{d}:=V_{d}^{\top}V_{d}$. In Proposition 3.1 we proved that $W_{d}$ is a diagonal invertible matrix, whenever the segments are unisolvent for histopolation. Hence, by left-multiplying for $W_{d}^{-1}$ and right-multiplying by $V_{d}^{-1}$, one gets
(14) $V_{d}^{-1}=W_{d}^{-1}V_{d}^{\top}.$
The elements for $V_{d}^{\top}$ are given in (9), while those of $W_{d}^{-1}$ are immediately retrieved by (10). Now, in general one has:
$\left[V_{d}^{-1}\right]_{i,j}$ $=\left[W_{d}^{-1}V_{d}^{\top}\right]_{i,j}=\sum_{k=1}^{d}\left[W_{d}^{-1}\right]_{i,k}\left[V_{d}^{\top}\right]_{k,j}=\sum_{k=1}^{d}\left[W_{d}^{-1}\right]_{i,k}\left[V_{d}\right]_{j,k}.$
Since $W_{d}^{-1}$ is diagonal, plugging in the explicit expressions of the terms computed in (10), one has
$\left[V_{d}^{-1}\right]_{i,j}=\left[W_{d}^{-1}\right]_{i,i}\left[V_{d}\right]_{i,j}=\frac{i^{2}}{2d\sin^{2}(i\rho)}\cdot\frac{2\sin(i\tau_{j})\sin(i\rho)}{i}=\frac{i\sin(i\tau_{j})}{d\sin(i\rho)}.$
3.2 A formula for the conditioning of Chebyshev segments
An immediate consequence of Proposition 3.1 is that the singular values of $V_{d}$ are the diagonal entries of $V_{d}^{\top}V_{d}$. A direct inspection on (10) also shows that the diagonal elements of $V_{d}^{\top}V_{d}$ are decreasingly ordered as, for $i>j$,
$\left[V_{d}^{\top}V_{d}\right]_{i,i}=\frac{2d}{i^{2}}\sin^{2}(i\rho)\geq\frac{2d}{j^{2}}\sin^{2}(j\rho)=\left[V_{d}^{\top}V_{d}\right]_{j,j}.$
This gives the simplified formula
(15) $\kappa_{2}\left(V_{d}\right)=\frac{\sigma_{1}(V_{d})}{\sigma_{d}(V_{d})}=\sqrt{\frac{\lambda_{1}(V_{d}^{\top}V_{d})}{\lambda_{d}(V_{d}^{\top}V_{d})}}=\sqrt{\frac{[V_{d}^{\top}V_{d}]_{1,1}}{[V_{d}^{\top}V_{d}]_{d,d}}}=\sqrt{\frac{d^{2}\sin^{2}(\rho)}{\sin^{2}(d\rho)}}=\left|\frac{d\sin(\rho)}{\sin(d\rho)}\right|$
for the conditioning of $V_{d}$, that hence depends on $\rho$. It is however quite natural to inversely relate $\rho$ and $d$, as done in [4], for instance in such a way that the arcs overlap in their endpoints, to obtain segments that belong to both classes (C1) and (C2). More generally, one has the following result.
Theorem 3.8.
Let $\rho=\frac{a}{d}$. Then,
$\lim_{d\to\infty}\kappa_{2}\left(V_{d}\right)=\left|\frac{a}{\sin(a)}\right|.$
In particular, for Chebyshev segments belonging to the intersection of classes (C1) and (C2), the conditioning $\kappa_{2}(V_{d})$ is monotonically increasing towards
$\lim_{d\to\infty}\kappa_{2}\left(V_{d}\right)=\pi/2.$
Proof 3.9.
Plugging the condition $\rho=\frac{a}{d}$ into (15), one gets that
$\kappa_{2}\left(V_{d}\right)=\left|\frac{d\sin(\rho)}{\sin(d\rho)}\right|=\left|\frac{d\sin(\frac{a}{d})}{\sin(d\frac{a}{d})}\right|=\frac{1}{|\sin(a)|}\left|d\sin\left(\frac{a}{d}\right)\right|.$
By considering the limit $d\to\infty$, and setting $t=a/d$, one has
$\lim_{d\to\infty}\kappa_{2}\left(V_{d}\right)=\lim_{d\to\infty}\frac{1}{|\sin(a)|}\left|d\sin\left(\frac{a}{d}\right)\right|=\lim_{t\to 0}\frac{|a|}{|\sin(a)|}\left|\frac{\sin\left(t\right)}{t}\right|=\frac{|a|}{|\sin(a)|}.$
The specification for $a=\pi/2$ is immediately obtained by substituting the value. To show the monotonicity of the conditioning, we shall prove that
$\frac{(d+1)\sin\left(\frac{\pi}{2(d+1)}\right)}{\sin\left((d+1)\frac{\pi}{2(d+1)}\right)}=\frac{(d+1)\sin\left(\frac{\pi}{2(d+1)}\right)}{\sin\left(\frac{\pi}{2}\right)}\geq\frac{d\sin\left(\frac{\pi}{2d}\right)}{\sin\left(\frac{\pi}{2}\right)}=\frac{d\sin\left(\frac{\pi}{2d}\right)}{\sin\left(d\frac{\pi}{2d}\right)}$
for each $d\geq 1$. This can be done by showing that the function $f(x):=x\sin(\pi/(2x))$ is increasing for $x>1$, i.e., that $f^{\prime}(x)>0$ for $x>1$. After differentiation, making the change of variable $t=\pi/(2x)$, it is readily seen that this is equivalent to proving that $g(t):=\tan(t)-t>0$ for $t\in(0,\pi/2)$. The first derivative of this function is $g^{\prime}(t)=\tan^{2}(t)$, which is clearly defined and positive in such an interval, and since $\lim_{t\to 0^{+}}g(t)=0^{+}$, this concludes the proof.
A numerical verification of the case $a=\pi/2$ of Theorem 3.8 is provided in Figure 7.
Figure 7: Bounded conditioning of the Vandermonde matrix in the Chebyshev basis for the class (C2). The parameter here is $a=\pi/2$, so the resulting segments also belong to the class (C1).
Let us now remark an interesting aspect. The selection of the basis is crucial to avoid an exponential conditioning of the Vandermonde matrix. Nevertheless, without an appropriate selection of histopolation segments, Chebyshev polynomials are not sufficient to control the conditioning of the Vandermonde matrix. This is depicted in the left hand panel of Figure 8, where the conditioning of the Chebyshev basis (blue solid line) on equidisitributed segments of class (C1) is shown. Although there is a clear improvement with respect to the monomial basis (red dashed line), the exponential growth for the conditioning is not avoided. On the other hand, the choice of the family of segments is of interest as well: the right hand panel of Figure 8 shows the conditioning of the monomial basis (blue solid line) for segments in the class (C2), which is smaller than that obtained for segments in class (C1) for the corresponding polynomial degrees (again, red dashed line). For comparison, equidistributed segments in the class (C1) on I=[1,1]I=[-1,1] were considered.
Figure 8: Left: conditioning of the Chebyshev basis for segments in the class (C1). Right: conditioning of the monomial basis for segments in class (C2), compared with segments in class (C1).
3.3 Conditioning with respect to the Frobenius norm
Proposition 3.1 and subsequent computations allow to deduce results beyond the conditioning in the Euclidean norm. Let us recall that, for a matrix $A\in\mathbb{R}^{m\times n}$, its Forbenius norm is the quantity
$\|A\|_{F}:=\sqrt{\sum_{i=1}^{m}\sum_{i=1}^{n}|A_{i,j}|^{2}}$
and the conditioning in the Frobenius norm is thus
(16) $\kappa_{F}(A):=\|A\|_{F}\|A^{-1}\|_{F}.$
The quantity $\kappa_{F}(A)$ reflects some geometrical features of the matrix $A$ [7], and in the context of nodal interpolation with Chebyshev nodes it has been studied in [9]. For large $d$, we are in a position to provide an asymptotic trend for $\kappa_{F}(V_{d})$. To this end, we shall need the integral sine function
$\mathrm{Si}(z):=\int_{0}^{z}\frac{\sin(t)}{t}\mathrm{d}t.$
Lemma 3.10.
One has
(17) $\lim_{d\to\infty}\|V_{d}\|_{F}=\sqrt{2a\mathrm{Si}(2a)+\cos(2a)-1}.$
In particular, if $a=\pi/2$, one has $\lim_{d\to\infty}\|V_{d}\|_{F}\leq 2$.
Proof 3.11.
First of all, recall that $\|V_{d}\|_{F}^{2}=\mathrm{Tr}\left(V_{d}^{\top}V_{d}\right)$. Hence, by Proposition 3.1, we get
$\|V_{d}\|_{F}^{2}=\mathrm{Tr}\left(V_{d}^{\top}V_{d}\right)=\sum_{i=1}^{d}\frac{2d}{i^{2}}\sin^{2}(i\rho)=\sum_{i=1}^{d}\frac{2d}{i^{2}}\sin^{2}\left(i\frac{a}{d}\right).$
The terms of this sum are positive and decreasing to 0, and the sum converges due to the presence of the term $i^{2}$. By the definition of Riemann integral, setting $\Delta x:=1/d$ and $x_{i}:=i/d$ and passing to the limit for $d\to\infty$, one has
$\lim_{d\to+\infty}\sum_{i=1}^{d}\frac{2d}{i^{2}}\sin^{2}\left(a\frac{i}{d}\right)=\lim_{d\to+\infty}\sum_{i=1}^{d}\frac{1}{x_{i}^{2}}\sin^{2}\left(ax_{i}\right)\Delta x=2\int_{0}^{1}\frac{\sin^{2}\left(ax\right)}{x^{2}}\mathrm{d}x.$
This integral can be expressed using the sine integral function $\mathrm{Si}(z)$, and after the immediate change of variable $t=ax$ and applying integration by parts, one obtains
$2\int_{0}^{1}\frac{\sin^{2}\left(ax\right)}{x^{2}}\mathrm{d}x=2a\int_{0}^{a}\frac{\sin^{2}(t)}{t^{2}}\mathrm{d}t=2a\mathrm{Si}(2a)+\cos(2a)-1.$
Taking the square root we get the claim, and substituting $a=\pi/2$ we immediately obtain the respective bound, as $\mathrm{Si}(\pi)\approx 1.85$.
Lemma 3.10 gives an exact formula for the first term of the right hand side of (16). Althouhgh such a neat result cannot be replicated for $\|V_{d}^{-1}\|_{F}$, we can at least estimate its trend as $d$ grows large.
Lemma 3.12.
Let $0<a\leq\pi/2$. One has
$\left\|V_{d}^{-1}\right\|_{F}\leq\frac{d}{\sqrt{2}a}.$
In particular, for $a=\pi/2$, one has $\|V_{d}^{-1}\|_{F}\leq d\frac{\sqrt{2}}{\pi}$.
Proof 3.13.
Again, let us observe that $\|V_{d}^{-1}\|^{2}_{F}=\mathrm{Tr}\left((V_{d}^{-1})^{\top}V_{d}^{-1}\right)=\mathrm{Tr}\left(V_{d}^{-\top}V_{d}^{-1}\right)$. Recalling formula (14), which states that $V_{d}^{-1}=W_{d}^{-1}V_{d}^{\top}$, being $W_{d}$ the diagonal matrix $W_{d}=V_{d}^{\top}V_{d}$, exploiting the cyclic property of the trace we get
$\left\|V_{d}^{-1}\right\|^{2}_{F}$ $=\mathrm{Tr}\left(V_{d}^{-\top}V_{d}^{-1}\right)=\mathrm{Tr}\left(V_{d}^{-\top}W_{d}^{-1}V_{d}^{\top}\right)=\mathrm{Tr}\left(V_{d}^{\top}V_{d}^{-\top}W_{d}^{-1}\right)=\mathrm{Tr}\left(W_{d}^{-1}\right).$
By Proposition 3.1, $W_{d}^{-1}$ is a diagonal matrix whose diagonal entries are the reciprocal of the diagonal entries of $V_{d}^{\top}V_{d}$, that is,
$\left[W_{d}^{-1}\right]_{i,i}=\frac{i^{2}}{2d\sin^{2}\left(\frac{a}{d}\right)}.$
Now, observe that, for $0\leq t\leq\pi/2$ one has $\sin(t)\geq t/2$. Since $a\leq\pi/2$ and $i/d\leq 1$, we may bound
$\mathrm{Tr}\left(W_{d}^{-1}\right)=\sum_{i=1}^{d}\frac{i^{2}}{2d\sin^{2}(a\frac{i}{d})}\leq\sum_{i=1}^{d}\frac{i^{2}}{2d(a\frac{i}{d})^{2}}=\sum_{i=1}^{d}\frac{d}{2a^{2}}=\frac{d^{2}}{2a^{2}}.$
By taking the square root of this quantity, we immediately get that $\|V_{d}^{-1}\|_{F}\leq\frac{d}{\sqrt{2}a}$.
We are in a position to prove the main result of the section.
Theorem 3.14.
The conditioning $\kappa_{F}(V_{d})$ grows asymptotically linearly with $d$.
Proof 3.15.
Lemma 3.10 gives an equality for the term $\|V_{d}\|_{F}$, while Lemma 3.12 gives an upper bound for $\|V_{d}^{-1}\|_{F}$. Since $\kappa_{F}(V_{d})=\|V_{d}\|_{F}\|V_{d}^{-1}\|_{F}$, the linear upper bound is established.
To obtain the lower bound, it is sufficient to observe that
$\sum_{i=1}^{d}\frac{i^{2}}{2d\sin^{2}\left(\frac{i}{d}a\right)}\geq\frac{1}{2d}\sum_{i=1}^{d}i^{2}=\frac{1}{2}\frac{d(d+1)(2d+1)}{6}\geq\frac{2d^{3}}{12d}=\frac{d^{2}}{6}.$
The square root of this quantity offers thus a linear lower bound for $\|V_{d}^{-1}\|_{F}$, whence the claim.
The result of Theorem 3.14 is depicted, for the limiting case $a=\pi/2$, in Figure 9.
Figure 9: Linear conditioning in the Frobenius norm of the Vandermonde matrix in the Chebyshev basis for classes (C2).
Conclusions and further directions
In the present paper, we have studied the conditioning of Vandermonde matrices $V_{d}$ arising in histopolation. We focused on the Euclidean conditioning $\kappa_{2}(V_{d})$, and we were able to identify two antipodal behaviours: the first one is exponential with $d$, and is peculiar to the monomial basis. While this basis is somehow sensitive to the underlying collection of supports, the general trend is indeed exponential, and hence gives a neat example of unifying exponential ill-conditioning, in all the $\|\cdot\|_{l^{k}\rightarrow l^{p}}$ norms, $k,p\in[1,\infty]$ (the case $k=p=2$ being the standard Euclidean or spectral norm). On the other hand, we recognized that the use of the Chebyshev basis can mitigate this effect; in particular, in combination with a particular class of segments, the conditioning drops and the Vandermonde matrix becomes very well-conditioned. As a by-product, owing to a discrete orthogonality, we also proved that this latter situation is favourable also with respect to the conditioning in the Frobenius norm $\kappa_{F}(V_{d})$.
Although the conditioning with respect to the Euclidean norm is the most common in numerical analysis, it is important to point out that it is not the only one. For instance, the conditioning with respect to the $\infty$-norm is usually taken into account when dealing with Bernstein-Vandermonde collocation matrices, whose numerical features are not only of theoretical interest, but also have several applications [1]. Thus, a natural extension to this work consists in considering the Bernstein basis (fixing the reference interval to I=[1,1]I=[-1,1]) and analyzing the relative conditioning in the $\infty$-norm. A well-established and interesting starting point for a worthwhile comparison with the literature is represented by the work [10].